can you give me an example of how to solve this question. i am a parent of a 12 year old and didn't take algebra when i was in school. any help would be greatly appreciated!!
3x + 6 = 6

Answer by rfer(16322)

3x+6=6
subtract 6 from both sides
3x+6-6=6-6
3x=0
divide both sides by 3, by you can't divide 0
so the answer is
x=0/3
x=0
Does that help?
Bob